OpenStudy (anonymous):

Find all the variables. 2B + i+K = 8 2i -k=2 4B +6i=20

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

not helpful. I need to find x,y,z but that surely is not the answer

OpenStudy (anonymous):

you can get rid of the k in the second equation, write \(k=2i-2\) plug that in the the first equation get \[2b+i+2i-2=8\] or just \[2b+3i=10\]

OpenStudy (anonymous):

but in this case there isnt x,y,z just three different variables

OpenStudy (anonymous):

then solve \[2b+3i=10\\ 4b+6i=20\]you see that there are an infinite number of solutions because it s=is the same line

OpenStudy (anonymous):

??????????? so confused still

OpenStudy (anonymous):

mk guess i wont get help.